OpenStudy (anonymous):

let x=2sin(theta), -pi/2 < theta < pi/2. Simplify the expression x/square root of 4-x^2

OpenStudy (anonymous):

make the replacement just like you are asked write \[\frac{2\sin(\theta)}{\sqrt{4-(2\sin(\theta))^2}}\]

OpenStudy (anonymous):

then the denominator can be manipulated as follows \[\sqrt{4-(2\sin(\theta))^2}=\sqrt{4-4\sin^2(\theta)}\] \[=\sqrt{4(1-\sin^2(\theta))}=\sqrt{4\cos^2(\theta)}=2\cos(\theta)\]